# Cold-Start Recommendation with Knowledge-Guided Retrieval-Augmented Generation

### v1

@@ 4.2.2. Ablation Study (RQ2 & RQ3)
To assess the contribution of ColdRAG’s main modules, we perform an ablation study targeting RQ2 and RQ3. We evaluate three variants: (1) w/o G,M, a plain LLM baseline without knowledge grounding (G) or multi-hop reasoning (M)—relying solely on user history and randomly sampled candidates; (2) w/o M, a ColdRAG variant that includes the knowledge graph (G) for entity retrieval but removes the reasoning component, retrieving only items matched directly to the user history without multi-hop traversal (M); and (3) the full ColdRAG, which leverages both G and M to construct contextualized prompts through dynamic graph traversal.
@@
Figure 4. Box plot comparison of Recall@10 and NDCG@10 across five runs for zero-shot baselines on the Games dataset.
@@ 4.2.3. Analysis on Generation Stability (RQ4)
LLMs are known to exhibit generation inconsistency—producing varied outputs for the same input depending on minor prompt changes or decoding randomness (Elazar et al., 2021). This inconsistency can undermine reliability in recommendation settings, where stable outputs are crucial for user trust. To investigate ColdRAG’s robustness, we assess generation stability by measuring the variance in recommendation quality across five runs. Figure 4 displays the performance distributions of Recall@10 and NDCG@10 using box plots for each zero-shot model, offering insight into output consistency.
ColdRAG consistently exhibits both high performance and low variance across runs, highlighting its ability to produce stable recommendation outputs. In contrast, other zero-shot baselines show wider distributions, indicating greater sensitivity to randomness during generation. Notably, we exclude KALM4Rec for fair comparison, as it prompts the LLM to generate item identifiers instead of natural language-formed item titles, making its output format incompatible with our evaluation setup.
This stability can be attributed to ColdRAG’s use of structured, knowledge-grounded retrieval via multi-hop reasoning over a semantic graph. Rather than relying solely on raw prompt inputs, ColdRAG constructs contextual signals that are consistent and semantically relevant, enabling the LLM to make more reliable decisions. These findings suggest that ColdRAG not only outperforms other models in accuracy but also provides a more deterministic and controllable recommendation process, which is particularly important in production environments where reproducibility matters.
Figure 5. Out-of-domain generation rate comparison on the Games dataset.
@@ 4.2.4. Analysis on Hallucination (RQ5)
A persistent issue in LLM-based recommendation is hallucination—the generation of item titles that do not belong to the target dataset. Such out-of-domain outputs reduce the reliability and usability of the recommendations in real-world applications. To evaluate ColdRAG’s effectiveness in mitigating hallucination, we analyze the proportion of out-of-domain items generated on the Games dataset. For a fair comparison, all baselines—LLM, LLMRank(S), LLMRank(R), LLMRank(I), and TaxRec—are provided with explicitly constructed candidate lists in their prompts to restrict the generation scope.
Despite this constraint, these models still produce a substantial number of out-of-domain items, with hallucination rates ranging from 5.82% to 9.89%. In contrast, ColdRAG achieves a significantly lower hallucination rate of just 3.15%, as shown in Figure 5. This indicates that even with access to candidate lists, prompt-based control alone is insufficient to prevent hallucination. ColdRAG eliminates the need for such predefined candidate lists by leveraging a dynamically constructed knowledge graph for retrieval. Its LLM-guided multi-hop reasoning mechanism allows it to retrieve semantically relevant, in-domain content grounded in structured knowledge. This grounding substantially reduces hallucination and improves the controllability and trustworthiness of the generated recommendations.
@@ 5. Conclusion
In this work, we introduced ColdRAG, a retrieval-augmented generation framework designed to address the item cold-start problem by integrating LLM-based reasoning with a dynamically constructed knowledge graph. ColdRAG first enriches item metadata into coherent textual descriptions, then extracts entities and relations to build a knowledge graph. It retrieves relevant item candidates by reasoning over this graph, and finally generates recommendations by combining these candidates with contextually grounded information in a prompt. Unlike prior approaches that depend on supervised training or handcrafted prompts, ColdRAG grounds generation in structured external knowledge and enables flexible, LLM-guided reasoning to identify semantically relevant candidates. This allows the framework to handle sparse metadata, enhance recommendation quality, and reduce out-of-domain generation—all without additional model training. Extensive experiments across multiple domains demonstrate that ColdRAG consistently outperforms both full-train and zero-shot baselines in accuracy, stability, and output reliability. Ablation studies validate the effectiveness and complementarity of its two core modules: knowledge graph construction and multi-hop reasoning. By unifying structured retrieval and generation, ColdRAG offers a practical and reliable solution for cold-start recommendation in zero-shot manner.
Looking forward, ColdRAG opens up promising directions for future research. One exciting direction is the development of agentic recommendation systems where LLMs actively explore, query, and refine their retrieval over multiple steps to improve recommendation outcomes. Another direction is integrating collaborative signals—such as latent patterns from user-item interactions—into the graph to better model user preferences. Additionally, extracting and incorporating implicit knowledge from user histories may further enhance recommendation quality and personalization.
GenAI Usage Disclosure
During the preparation of this work, we used generative AI tools solely for writing-related tasks, such as correcting spelling and grammar, improving clarity, and fixing LaTeX formatting errors in tables and equations. All technical contributions, ideas, experiments, and analyses were solely developed and verified by the authors without any assistance from generative AI systems. No generative AI tools were used for coding, data processing, or result generation. All outputs produced with generative AI assistance were carefully reviewed and edited by the authors to ensure accuracy and adherence to academic standards. This use complies with the ACM policy on authorship and the use of generative AI tools in scholarly publishing.

### v2

@@ 4.2.2. Ablation Study (RQ2 & RQ3)
To assess the contribution of ColdRAG’s main modules, we perform an ablation study, evaluating three variants: (1) w/o G,M, a plain LLM baseline without knowledge grounding (G) or multi-hop reasoning (M)—relying solely on user history and randomly sampled candidates; (2) w/o M, a ColdRAG variant that includes the knowledge graph (G) for entity retrieval but removes the reasoning component, retrieving only items matched directly to the user history without multi-hop traversal (M); and (3) the full ColdRAG, which leverages both G and M to construct prompts through dynamic graph traversal.
@@
Figure 4. (a) Box plot comparison of Recall@10 across five runs for zero-shot baselines on the Games dataset. (b) Out-of-domain generation rate comparison on the Games dataset.
@@ 4.2.3. Analysis on Generation Stability (RQ4)
LLMs are known to exhibit generation inconsistency—producing varied outputs for the same input depending on minor prompt changes or decoding randomness (Elazar et al., 2021). This inconsistency can undermine reliability in recommendation settings, where stable outputs are crucial for user trust. To investigate ColdRAG’s robustness, we assess generation stability by measuring the variance in recommendation quality across five runs. Figure 4 (a) displays the performance distributions of Recall@10 on the Games dataset using box plots for each zero-shot model, offering insight into output consistency.
ColdRAG consistently exhibits both high performance and low variance across runs, highlighting its ability to produce stable recommendation outputs. In contrast, other zero-shot baselines show wider distributions, indicating greater sensitivity to randomness during generation. Notably, we exclude KALM4Rec for fair comparison, as it prompts the LLM to generate item identifiers instead of natural language-formed item titles, making its output format incompatible with our evaluation setup. This stability can be attributed to ColdRAG’s use of structured, knowledge-grounded retrieval via multi-hop reasoning over a semantic graph. Rather than relying solely on raw prompt inputs, ColdRAG constructs contextual signals that are consistent and semantically relevant, enabling the LLM to make more reliable decisions. These findings suggest that ColdRAG not only outperforms other models in accuracy but also provides a more deterministic and controllable recommendation process, which is particularly important in production environments where reproducibility matters.
@@ 4.2.4. Analysis on Hallucination (RQ5)
A persistent issue in LLM-based recommendation is hallucination—the generation of invalid item titles that are absent from the target dataset. Such out-of-domain outputs reduce the reliability and usability of the recommendations in real-world applications. To evaluate ColdRAG’s effectiveness in mitigating hallucination, we analyze the proportion of out-of-domain items generated on the Games dataset, as shown in Figure 4. For a fair comparison, all baselines—LLM, LLMRank(S), LLMRank(R), LLMRank(I), and TaxRec—are provided with explicitly constructed candidate lists in prompts for controlled generation.
Despite this constraint, these models still produce a substantial number of out-of-domain items, with hallucination rates ranging from 5.82% to 9.89%. In contrast, ColdRAG achieves a significantly lower hallucination rate of just 3.15%. This indicates that even with access to candidate lists, prompt-based control alone is insufficient to prevent hallucination. ColdRAG eliminates the need for such predefined candidate lists by leveraging a dynamically constructed knowledge graph for retrieval. Its LLM-guided multi-hop reasoning mechanism allows it to retrieve semantically relevant, in-domain content grounded in structured knowledge. This grounding substantially reduces hallucination and improves the controllability and trustworthiness of the generated recommendations.
Figure 5. (a) Pie chart of the distribution of entity (node) types in the KG. (b) Heatmap of the number of relations (edges) between the entity types.
4.2.5. Structure of Knowledge Graph
To provide better understanding about the structure of the generated KG constructed in Knowledge Graph Construction, we analyzed the composition of its nodes and edges in the Games domain. Figure 5 (a) illustrates the distribution of node types. The KG is predominantly composed of item nodes, which represent half of all entities (50%), and feature nodes, making up another significant portion (26%). Other entity types such as target user (9%), etc (6%), setting (5%), and genre (4%) constitute the remainder, providing rich, multi-faceted context for the items. This distribution highlights that the KG is naturally centered around the items themselves, as all other non-item entities are components generated to describe and explain the items. The feature category is the second most abundant because it serves as a broad, versatile concept that can capture a wider array of attributes for an item.
The heatmap in Figure 5 (b) details the connectivity between these different node types, showing the number of edges. A key observation is the dense connection between item and feature nodes, with 15,102 edges, confirming that items are primarily defined by a rich set of features within the graph. Similarly, the substantial number of edges between item and target user nodes (4,928) indicates that the graph effectively captures relationships relevant to user preferences. The connections between items and other concepts like genre (2,542) and setting (2,735) further enrich the semantic network, enabling more nuanced, multi-hop reasoning during the recommendation process. This interconnected structure demonstrates that the KG successfully models the complex relationships between items, their attributes, and user contexts, forming a solid foundation for the retrieval and reasoning stages of ColdRAG.
@@ 5. Conclusion
In this work, we introduced ColdRAG, a retrieval-augmented generation framework that addresses the item cold-start problem by integrating LLM-based reasoning with a dynamically constructed knowledge graph. Unlike typical graph-RAG systems that rely on simple text similarity, ColdRAG employs LLM-guided, multi-hop reasoning to retrieve evidence tailored to user preferences. This targeted approach allows the framework to handle sparse metadata, enhance recommendation quality, and reduce out-of-domain generation without requiring task-specific training. Extensive experiments validate that ColdRAG consistently outperforms state-of-the-art baselines in accuracy, stability, and reliability, offering a practical zero-shot solution for cold-start recommendation.